# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ 4.2 Lagrangian Data Assimilation of a High-Dimensional Flow Field
Lagrangian data assimilation (LDA) is a specialized method that utilizes the trajectories of moving tracers (e.g., drifters or floaters) as observations used to recover the unknown flow field driving their motion [42, 47, 119, 120, 121, 122]. Unlike Eulerian observations, which are fixed at specific locations, Lagrangian drifters track the movement of fluid parcels over time [123, 124, 125, 126]. This approach is particularly valuable for recovering ocean states in the mid-latitude using floats or in marginal ice zones (MIZ) using sea ice floes as Lagrangian tracers [53, 127, 128]. However, due to the high dimensionality of the state space of the flow field, which is characterized by a large number of spectral modes or high-resolution mesh grids, running a forward-backward offline smoother for state estimation becomes computationally infeasible. Consequently, the adaptive online smoother is essential in practical applications.
@@
${\,\rm d}\mathbf{v}_{\ell}(t)$ $=\beta\left(\mathbf{u}(t,\mathbf{x}_{\ell})-\mathbf{v}_{\ell}(t)\right){\,\rm d}t+\boldsymbol{\Sigma}^{\mathbf{v}_{\ell}}{\,\rm d}\mathbf{W}_{\mathbf{v}_{\ell}}(t),$ (4.4)
${\,\rm d}\hat{u}_{\mathbf{k}}(t)$ $=(-d_{\mathbf{k}}\hat{u}_{\mathbf{k}}(t)+f_{\mathbf{k}}(t)){\,\rm d}t+\sigma_{\mathbf{k}}{\,\rm d}W_{\mathbf{k}}(t),\qquad\text{with~{}}\mathbf{u}(t,\mathbf{x}_{\ell}(t))=\sum_{\mathbf{k}\in\mathscr{K}}\hat{u}_{\mathbf{k}}(t)e^{i\mathbf{k}\cdot\mathbf{x}_{\ell}(t)}\mathbf{r}_{\mathbf{k}},$ (4.5)
@@
where $\mathbf{x}_{\ell}$ and $\mathbf{v}_{\ell}$ denote the location and velocity of the $\ell$-th tracer, respectively, with $\ell=1,\ldots,L$, while $\mathbf{u}(t,\mathbf{x})$ is the ocean velocity field expressed through a spectral decomposition in its geostrophically balanced or potential vortical incompressible modes, using a Fourier basis in the doubly-periodic spatial domain $[-\pi,\pi]$ [129]. The governing equations of the Fourier coefficients $\hat{u}_{\mathbf{k}}$ are expressed through a set of linear stochastic processes, where $\mathbf{k}\in\mathscr{K}$ denotes the Fourier wavenumber with $\mathscr{K}=[-K,K]^{2}\cap\mathbb{Z}^{2}$ for $K\geq 1$ denoting the two-dimensional discrete lattice collection of Fourier wavenumbers. Note that the stochastic noise mimics the turbulent effects of nonlinearity present in many practical systems, such as the quasi-geostrophic (QG) ocean model. This approximation is justified in many applications [42, 47]. Since the primary goal is to illustrate effectiveness of the online smoother, this simplification is employed in the study here. By treating all the $\mathbf{x}_{\ell}$ as the observational variables and all the $\mathbf{v}_{\ell}$ and $\hat{u}_{\mathbf{k}}$ as the variables for state estimation, the LDA system in (4.3)–(4.5) belongs to the CGNS family. The parameter values in this simulation are the following. The wavenumber bound is $K=2$ such that there are in total 25 modes after excluding the zeroth mode which corresponds to a time-varying random background mean sweep or background velocity field. The total simulation time is $T=5$ with the numerical integration time step (also the observational time frequency) being $\Delta t=0.005$. The total number of tracers is $L=18$. The model parameters are $f_{\mathbf{k}}(t)=0.15e^{\displaystyle 5\pi it}$, $\boldsymbol{\Sigma}^{\mathbf{x}_{\ell}}=0.005\pi\mathbf{I}_{2\times 2}$, $\beta=1$, $\boldsymbol{\Sigma}^{\mathbf{v}_{\ell}}=0.1\mathbf{I}_{2\times 2}$. The damping and noise coefficients $d_{\mathbf{k}}=d_{-\mathbf{k}}$ and $\sigma_{\mathbf{k}}=\sigma_{-\mathbf{k}}$ are randomly drawn from $\mathcal{U}([0.5,1.5])$ and $\mathcal{U}([0.15,0.25])$, respectively, with the conjugacy condition establishing the reality of the underlying velocity field [42, 47].
@@
Figure 4.3 demonstrates the posterior state estimates using the filter (Theorem 2.2) and smoother (Theorem 2.3) in recovering the underlying ocean flow field and the tracer velocity vectors. Note that the amplitudes for the ocean flow field (in blue) and tracer velocity (in the hot colormap) are not shown on the same scale. Panels (a)–-(c) compare the spatial recovery of the ocean state and tracer velocity vectors at $t=2$. Both methods effectively reconstruct the tracer movement in direction and magnitude, with the smoother being slightly more effective. As for the flow velocity field, the smoother exhibits a much higher skill in its recovery compared to the filter solution. This is further illustrated in panels (d)-–(e), where we compare the true time series with the filter and smoother posterior mean states over the observational period for the real part of the ocean Fourier mode $\mathbf{k}=(2,-1)^{\mathtt{T}}$ of the ocean, and the zonal velocity of tracer #1. We can see that the smoother follows true values more closely and with less uncertainty for both cases. This is observed uniformly among all Fourier modes and all tracer velocity vectors.
@@
Figure 4.3: Panel (a): True state of the underlying flow field $t=2$, where the colormap shows the amplitude and the quiver plot represents the velocity field. It is superimposed by the location of the tracers and their velocity vectors. The magnitudes of the quiver plot for the ocean’s velocity field and velocity vectors for the tracers are on distinct scales. Panels (b)–(c): Similar to (a), but corresponding to the estimated state from the posterior mean, calculated through the filter and smoother, respectively. Panels (d)–(e): Comparison between the true time series (in blue) and the posterior mean time series of the filter (in green) and smoother (in red) solutions for the real part of the ocean Fourier mode $\mathbf{k}=(2,-1)^{\mathtt{T}}$ and zonal velocity of tracer #1 (labeled in panels (a)–(c)).
@@
Figure 4.4 compares the computational time (in seconds) and storage requirements (in gigabytes) for the fixed- and adaptive-lag strategies of the online smoother. Additionally, we present the normalized root-mean-square error (NRMSE) of the posterior mean time series to assess the pathwise accuracy of the methods. The normalized RMSE is computed by dividing the RMSE (calculated in the temporal direction) by the standard deviation of the true signal [130, 131, 132], such that the NRMSE indicates an error level comparable to the equilibrium variability. This metric is evaluated as functions of the fixed lag and tolerance parameter $\delta$ and showed on a logarithmically scaled x-axis, where for the latter, a maximum lag bound of $b=300$ is applied throughout and the sequence of entropies is used to define the adaptive lags (i.e., (3.33)). Panels (a) and (b) display fixed-lag results, while panels (d) and (e) display the adaptive-lag strategies. To provide robust results, computational times in panels (b) and (e) are averaged over multiple runs to reduce external fluctuations. For the fixed-lag smoother, both storage and time increase linearly as functions of the fixed-lag, while the NRMSE converges exponentially from the filter solution (zero lag) to the one for the smoother solution (maximum lag), indicating the sufficiency of using an overall short lag to significantly save on computational storage. On the other hand, the storage for the adaptive-lag smoother remains constant with respect to $\delta$ since a fixed upper bound $b$ is being used in all cases, while computational time increases exponentially as $\delta$ approaches zero. The NRMSE for the adaptive-lag method now decreases linearly with tolerance, demonstrating a sharper trade-off between computational resources and error compared to the fixed-lag approach. Panels (c) and (f) further illustrate the recovery skill of these approaches through the temporal average of the information gain for the fixed- and adaptive-lag online smoother strategies beyond the filter solution as functions of the fixed-lag and tolerance parameter, respectively. Both are plotted on an x-axis that is scaled logarithmically and are calculated using the signal-dispersion formula of the relative entropy in (3.26) since both distributions are Gaussian at each time instant. For reference, we show both the signal and dispersion parts, as well as the total information gain. For the fixed-lag smoother, since the zero fixed-lag online smoother coincides with the filter solution, the information gain there is zero, but as the fixed-lag increases, it rapidly and exponentially converges to the information gain value corresponding to the full-backward offline smoother beyond the filter estimates (see Section 3.3). On the other hand, similarly to the results from panels (d) and (e), we have that the temporally-averaged information gain beyond the filter solution is a linear function in $\delta$, where for large values of $\delta$ this is close to zero and linearly approaches the information gain value of the full-backward offline smoother beyond the filter estimates as $\delta$ approaches zero. In both cases, these behaviors are demonstrated in both the signal and dispersion parts and, as such, in the total information gain as well.
@@
Figure 4.4: Panels (a)–(b) & (d)–(e): Analysis of the trade-off between the computational complexity, measured in computational time (in seconds), and required storage capacity (in gigabytes), and lower-order pathwise error statistic of the NRMSE between the posterior mean time series and true signal, for the fixed-lag (panels (a)–(b)) and adaptive-lag online smoother (panels (d)–(e)). All are plotted on a semi-log x-axis, and the x-axis for the adaptive-lag online smoother is reversed for ease of comparison with the corresponding fixed-lag strategy plots. Panels (c) & (f): Temporal average of the information gain for the fixed- (panel (c)) and adaptive-lag (panel (f)) online smoother beyond the filter solution for various values of the fixed lag and tolerance parameter, respectively. Both the signal (in green) and dispersion (in blue) parts of the total information gain (in red) are showed. Their x-axes are logarithmically scaled, and for the adaptive-lag online smoother, it has also reversed direction.
@@
In Figure 4.5, we demonstrate the contrasting behavior of the adaptive-lag online smoother in the LDA problem compared to the dyad-interaction model with intermittent instabilities from Section 4.1. First, in panel (a), similar to panel (h) of Figure 4.2, the information gain criterion used to calculate the adaptive lag values in (3.33) is shown at different time instants for $R_{n}\leq j\leq n-1$ (with the x-axis is measured in simulation time units). In panel (b), same as in panel (i) of Figure 4.2, the spectral radii of the update tensors $\mathbf{D}^{j,n-2}$ are plotted at the same time instants as the sequences in panel (a), for $10\leq j\leq n-$ 1, where for both the y-axis is on a log scale. It is important to observe here the differences and similarities between these plots and the respective ones from Figure 4.2, for the highly intermittent and unstable dyad-interaction model defined in (4.1)–(4.2). In both case studies, these sequences illustrate an exponential behavior in $j$ for all values of $n$. However, for the LDA problem, we can see a uniform rate of increase for these metrics across the board as $j$ increases, due to the parallel lines of constant slope that emerge in both panels. Recall that, in the dyad-interaction numerical experiment, this rate would change depending on whether a rare or extreme event was forming in the observable signal; it could even exhibit regime switches depending on the emergence of intermittent instabilities. This reveals the fact that for the LDA problem of recovering the flow field’s and tracers’ velocities, the observations from the tracers have a uniform impact on the online smoother estimate regardless of when they are observed, which is a result of the stability that the model in (4.3)–(4.5) has, as well as of the fact that the equilibrium distribution of the tracers is nearly uniform due to the incompressibility of the underlying ocean velocity field [42]. This contrasts with the dyad-interaction model, which displays intermittent instabilities due to the latent dynamics acting as anti-damping in the observable ones on rare occasions. Despite this discrepancy, it is noted that the spectral radii of the update tensors $\mathbf{D}^{j,n-2}$, which are high-dimensional matrices in this case, remain below $1$ throughout time and for all values of $n$. This uniform influence region among the observations is also observed in panel (c). Similar to panel (g) in Figure 4.2, it illustrates the adaptive lag values of the online smoother with hyperparameters $b=100$ and $\delta=0.05$ as a function of time, where the adaptive lag is measured in simulation time units (i.e., $L_{n}\Delta t$) and calculated via the sequence of entropies through (3.33). Unlike the adaptive lag values generated from the dyad model in Section 4.1, which highly depend on whether an extreme event is being observed or if we are currently at a period of dormancy, the adaptive lag value here meanders around a constant (about $0.3$ in simulation time units), via uniform, in amplitude, oscillations (after, of course, $n-1\geq b+1$, since for $n-1\leq b$ the adaptive lag value is simply equal to $n-1$ due to (3.32)), and does so steadily in time which indicates the absence of intermittency in the system and the uniform impact each observation has on the online smoother solution.
Figure 4.5: Panel (a): The information gain criterion used to define the adaptive lags through (3.33), as a function of $R_{n}\leq j\leq n-1$, for various values of $n$ (i.e., time instants in $[0,5]$). The x-axis is in simulation time units. The threshold value of interest in this study, $\delta=0.05$, is also plotted via a black dashed line for reference. Panel (b): Spectral radii of the update tensors $D^{j,n-2}$ as a function of $10\leq j\leq n-$ 1 (x-axis is in simulation time units), for the same time instants as the ones displayed in panel (a). Panels (a) and (b) share the same legend. Panel (c): The adaptive lag values (in time units, i.e., $L_{n}\Delta t$) generated for the adaptive-lag online smoother with hyperparameters $b=100$ and $\delta=0.05$ when using (3.33) for their calculation.

### v2

@@ 4.2 Lagrangian Data Assimilation of a High-Dimensional Flow Field
Lagrangian data assimilation (LDA) is a specialized method that utilizes the trajectories of moving tracers (e.g., drifters or floaters) as observations used to recover the unknown flow field driving their motion [41, 42, 92, 93, 94]. Unlike Eulerian observations, which are fixed at specific locations, Lagrangian drifters track the movement of fluid parcels over time [95, 96, 97]. This approach is particularly valuable for recovering ocean states in the mid-latitude using floats or in marginal ice zones using sea ice floes as Lagrangian tracers [47, 98, 99]. However, due to the high dimensionality of the state space of the flow field, which is characterized by a large number of spectral modes or high-resolution mesh grids, running a forward-backward offline smoother for state estimation becomes computationally infeasible. Consequently, the adaptive online smoother is essential in practical applications.
@@
${\,\rm d}\mathbf{v}_{\ell}(t)$ $=\beta\big(\mathbf{u}(t,\mathbf{x}_{\ell})-\mathbf{v}_{\ell}(t)\big){\,\rm d}t+\boldsymbol{\Sigma}^{\mathbf{v}_{\ell}}{\,\rm d}\mathbf{W}_{\mathbf{v}_{\ell}}(t),$ (4.5)
${\,\rm d}\hat{u}_{\mathbf{k}}(t)$ $=\big(-d_{\mathbf{k}}\hat{u}_{\mathbf{k}}(t)+f_{\mathbf{k}}(t)\big){\,\rm d}t+\sigma_{\mathbf{k}}{\,\rm d}W_{\mathbf{k}}(t),\quad\text{with~}\mathbf{u}(t,\mathbf{x}_{\ell}(t))=\sum_{\mathbf{k}\in\mathscr{K}}\hat{u}_{\mathbf{k}}(t)e^{i\mathbf{k}\cdot\mathbf{x}_{\ell}(t)}\mathbf{r}_{\mathbf{k}},$ (4.6)
@@
where $\mathbf{x}_{\ell}$ and $\mathbf{v}_{\ell}$ denote the location and velocity of the $\ell$-th tracer, respectively, with $\ell=1,\ldots,L$, while $\mathbf{u}(t,\mathbf{x})$ is the ocean velocity field expressed through a spectral decomposition in its geostrophically balanced or potential vortical incompressible modes, using a Fourier basis in the doubly-periodic spatial domain $\pi,\pi] [$100]. The governing equations of the Fourier coefficients $\hat{u}_{\mathbf{k}}$ are expressed through a set of linear stochastic processes, where $\mathbf{k}\in\mathscr{K}$ denotes the Fourier wavenumber with $\mathscr{K}=[-K,K]^{2}\cap\mathbb{Z}^{2}$ for $K\geq 1$ denoting the two-dimensional discrete lattice collection of Fourier wavenumbers. Note that the stochastic noise mimics the turbulent effects of nonlinearity present in many practical systems, such as the quasi-geostrophic (QG) ocean model. This approximation is justified in many applications [41, 42]. Since the primary goal is to illustrate the effectiveness of the online smoother, this simplification is employed in the study here. By treating all the $\mathbf{x}_{\ell}$ as the observational variables and all the $\mathbf{v}_{\ell}$ and $\hat{u}_{\mathbf{k}}$ as the variables for state estimation, the LDA system in (4.4)–(4.6) belongs to the CGNS family. The parameter values in this simulation are the following. The wavenumber bound is $K=2$, such that there are in total 25 modes after excluding the zeroth mode, which corresponds to a time-varying random background mean sweep, or background velocity field. The total simulation time is $T=5$ with the numerical integration time step (also the observational time frequency) being $\Delta t=0.005$. The total number of tracers is $L=18$. The model parameters are $f_{\mathbf{k}}(t)=0.15e^{\displaystyle 5\pi it}$, $\boldsymbol{\Sigma}^{\mathbf{x}_{\ell}}=0.005\pi\mathbf{I}_{2\times 2}$, $\beta=1$, $\boldsymbol{\Sigma}^{\mathbf{v}_{\ell}}=0.1\mathbf{I}_{2\times 2}$. The damping and noise coefficients $d_{\mathbf{k}}=d_{-\mathbf{k}}$ and $\sigma_{\mathbf{k}}=\sigma_{-\mathbf{k}}$ are randomly drawn from $\mathcal{U}\big([0.5,1.5]\big)$ and $\mathcal{U}\big([0.15,0.25]\big)$, respectively, with the conjugacy condition establishing the reality of the underlying velocity field [41, 42].
@@
Figure 4.3 demonstrates the posterior state estimates using the filter (Theorem 2.2) and smoother (Theorem 2.3) in recovering the underlying ocean flow field and the tracer velocity vectors. Note that the amplitudes for the ocean flow field (in blue) and tracer velocity (in the hot colormap) quiver plots are not shown on the same scale. Panels (a)–-(c) compare the spatial recovery of the ocean state and tracer velocity vectors at $t=2$. Both methods effectively reconstruct the tracer movement in direction and magnitude, with the smoother being slightly more effective. As for the flow velocity field, the smoother exhibits a much higher skill in its recovery compared to the filter solution. This is further illustrated in Panels (d)-–(e), where we compare the true time series with the filter and smoother posterior mean states over the observational period for the real part of the ocean Fourier mode $\mathbf{k}=(2,-1)^{\mathtt{T}}$ of the ocean, and the zonal velocity of tracer #1 (noted in Panels (a)–(c)). We can see that the smoother follows the truth more closely and with less uncertainty for both cases. This is observed uniformly among all Fourier modes and all tracer velocity vectors.
@@
Figure 4.3: Panel (a): True state of the underlying flow field at $t=2$, where the colormap shows the amplitude and the quiver plot represents the velocity field. It is superimposed by the location of the tracers and their velocity vectors. The magnitudes of the quiver plot for the ocean’s velocity field and velocity vectors for the tracers are on distinct scales. Panels (b)–(c): Similar to (a), but corresponding to the estimated state from the posterior mean, calculated through the filter and smoother, respectively. Panels (d)–(e): Comparison between the true time series (in blue) and the posterior mean time series of the filter (in green) and smoother (in red) solutions for the real part of the ocean Fourier mode $\mathbf{k}=(2,-1)^{\mathtt{T}}$ and zonal velocity of tracer #1 (labeled in Panels (a)–(c)).
@@
Figure 4.4 compares the computational time (in seconds) and storage requirements (in gigabytes) for the fixed- and adaptive-lag strategies of the online smoother. (For details on how the storage and time values in Panels (a), (b), (d), and (e) are calculated, see Appendix D.) Additionally, we present the normalized root-mean-square error (NRMSE) between the posterior mean time series and the truth as to assess the pathwise accuracy of these methods. The normalized RMSE is computed by dividing the RMSE (calculated in the temporal direction) by the standard deviation of the true signal [101, 102], such that the NRMSE indicates an error level comparable to the equilibrium variability. This metric is evaluated as a function of the fixed lag and tolerance parameter $\delta$ and showed on a logarithmically scaled x-axis, where for the latter a maximum lag bound of $b=300$ (corresponding to $b\Delta t=1.5$ simulation time units) is applied throughout, with the sequence of entropies used to define the adaptive lags (i.e., (3.32) and (3.31)). Panels (a) and (b) display the fixed-lag results, while Panels (d) and (e) display the adaptive-lag ones. To provide robust results, computational times in Panels (b) and (e) are averaged over multiple runs to reduce external fluctuations. For the fixed-lag smoother, both storage and time increase algebraically (e.g., linearly) with lag, while the NRMSE converges exponentially, at a significant rate, from the filter solution (zero lag) to the one offline smoother one (maximum lag), indicating the sufficiency of using an overall short lag to significantly save on computational storage for this specific problem instance. On the other hand, the storage needs of the adaptive-lag smoother remain constant with respect to $\delta$, since a fixed $b$ is being used in all cases, while computational time increases algebraically to logarithmically as $\delta$ approaches zero. The former result showcases how the adaptive-lag online smoother implicitly minimizes storage requirements when compared to the fixed-lag variant, since for $b$ fixed $\delta$ can decrease freely to improve the recovery skill, with storage needs remaining unaffected. Note how by using a $b\Delta t$ value close to the uniform impact region of the observations in this LDA problem for a target pathwise recovery skill (e.g., $b\Delta t\approx 0.35$ for a target NRMSE of around $0.35$, where we have diminishing returns with decreasing $\delta$; see Panel (a) in Figure 4.5), there is a potential for significant storage savings to be had in this approach. Of course, the computational time spent unavoidably increases with decreasing $\delta$, since more linear-search checks in (3.32) (or (3.30)) need to be carried out. In Panel (e) we also note at each data point on the time plot the percentage of time spent on the calculation of the adaptive lag (also averaged over many runs); in this calculation we include the time spent evaluating the update matrices $\mathbf{D}^{j,n}$ for $n$ and $R_{n}\leq j\leq n-1$, since they are required for the calculation of $L_{n}$ (see (3.28)–(3.29)). The NRMSE for the adaptive-lag method now decreases algebraically with $\delta$, demonstrating a sharper trade-off between computational resources and error compared to the fixed-lag approach.
Panels (c) and (f) further illustrate the recovery skill of these approaches through the temporal average of the information gain for the fixed- and adaptive-lag online smoother strategies beyond the filter solution as functions of the fixed lag and tolerance parameter, respectively. Both are plotted on an x-axis that is scaled logarithmically and are calculated using the signal–dispersion formula of the relative entropy in (3.25), since both distributions are Gaussian at each time instant. For reference, we show both the signal and dispersion parts, as well as the total information gain. For the fixed-lag smoother, since the zero-fixed-lag online smoother coincides with the filter solution, the information gain there is zero, but as the fixed lag increases, it rapidly and exponentially converges to the information gain value corresponding to the full-backward offline smoother beyond the filter estimates (see Section 3.3). On the other hand, similarly to the results from Panels (d) and (e), we have that the temporally-averaged information gain beyond the filter solution is a logarithmic to algebraic function in $\delta$, where for large values of $\delta$ this is close to zero and approaches the information gain value of the full-backward offline smoother beyond the filter estimates as $\delta$ approaches zero. In both cases, these behaviors are demonstrated in both the signal and dispersion parts and, as such, in the total information gain as well.
@@
Figure 4.4: Panels (a)–(b) & (d)–(e): Analysis of the trade-off between the computational complexity, measured in computational time (in seconds), and required storage capacity (in gigabytes), and lower-order pathwise error statistic of the NRMSE between the posterior mean time series and true signal, for the fixed-lag (Panels (a)–(b)) and adaptive-lag online smoother (Panels (d)–(e)). In Panel (e) we also show at each data point of the latter the percentage of time spent on the adaptive-lag calculation (including the calculation of the update matrices $\mathbf{D}^{j,n-2}$). Details on how the storage and time values are calculated are given in Appendix D. Panels (c) & (f): Temporal average of the information gain for the fixed- (Panel (c)) and adaptive-lag (Panel (f)) online smoother beyond the filter solution for various values of the fixed lag and tolerance parameter, respectively. Both the signal (in green) and dispersion (in blue) parts of the total information gain (in red) are shown. All panels are plotted on a semi-log x-axis, and the x-axes for the adaptive-lag online smoother results are reversed for ease of comparison with the corresponding fixed-lag strategy plots.
@@
Finally, in Figure 4.5, we demonstrate the contrasting behavior of the adaptive-lag online smoother in the LDA problem when compared to the dyad-interaction model with intermittent instabilities from Section 4.1. First, Panel (a), similar to Panel (g) in Figure 4.2, illustrates the adaptive lag values (in time units) of the online smoother with hyperparameters $b=100$ (corresponding to $b\Delta t=0.5$ simulation time units), $\delta=0.05$ and calculated via the sequence of entropies through (3.32) and (3.31), as a function of time in $[2,5]$. Unlike the adaptive lag values generated from the dyad model in Section 4.1, which highly depend on whether an extreme event is being observed or if we are currently at a period of dormancy, the adaptive lag value here meanders around a constant (about $0.3$ in simulation time units), via uniform, in amplitude, oscillations, and does so steadily in time which indicates the absence of intermittency in the system and the uniform impact each observation has on the online smoother solution. Similar to Panels (h) and (i) of Figure 4.2, Panels (b) and (c) depict the standardized information gain sequence (see (4.3)) and spectral radii of the update matrices $\mathbf{D}^{j,n-2}$ but for the LDA problem defined in this section, (4.4)–(4.6). Again, as with the results from the dyad-interaction model, the two metrics showcase significant temporal correlations, with their major discrepancies emerging in the differing scale and noisier behavior of the former. As before, this functions as numerical corroboration to the claim that our proposed information-theoretic approach to determining the observational impact region of the $n$-th observation onto the online discrete smoother update is dynamically- and statistically-consistent and effective in recovering the exact approach reflected by the spectrum of the update tensor $\mathbf{D}^{j,n-2}$, while being much more computationally efficient at the same time. But, it is also important to compare and contrast these results with those from Panels (g)–(i) of Figure 4.2. Foremost, the spectral radii of the update tensors $\mathbf{D}^{j,n-2}$, which are high-dimensional matrices in this case, remain below $1$ throughout time and for all values of $n$. Also, in both case studies, these sequences illustrate an exponential behavior in $j$ for all values of $n$. However, for the LDA problem, we can see a uniform rate of exponential decay as $j$ decreases from n1n-1, across all observations (i.e., $n$). Recall that, in the dyad-interaction numerical experiment, this rate would change depending on whether a rare or extreme event was forming in the observable signal; it could even exhibit regime switches depending on the emergence of intermittent instabilities. Therefore, for the LDA problem where we recover the flow field’s and tracers’ velocities, the observations from the tracers’ locations have a uniform impact on the online smoother estimate over time, which is a result of the stability that the model in (4.4)–(4.6) has, as well as of the fact that the equilibrium distribution of the tracers is nearly uniform due to the incompressibility of the underlying ocean velocity field [41]. This contrasts with the dyad-interaction model, which displays intermittent instabilities due to the latent dynamics acting as anti-damping in the observable ones on rare occasions.
Lastly, in Panel (d) of Figure 4.5, we showcase the standardized LSDev criterion, i.e., for $\sigma^{j,n}$ in (4.3), in the same manner as Panel (b). In this case study, where the absence of the intermittent instabilities allows for an almost consistent exponentially decaying behavior with a constant rate in $j$ for each $n$ for the adaptive-lag-defining criteria, Panels (b) and (d) numerically corroborate the note made in the second-to-last paragraph in Section 3.4.2; we can yield analogous adaptive lag values with either sequence (of relative entropies $\left\{\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)\right\}_{R_{n}\leq j\leq n-1}$ or of their LSDevs $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}$) by a simple adjustment of the $\delta$ tolerance value.
Figure 4.5: Panels (a)–(c): Similar to Panels (g)–(i) of Figure 4.2, respectively. Panel (d): Same as Panel (b) but for the standardized LSDev sequence (i.e., for $\sigma^{j,n}$ in (4.3)). All panels correspond to the LDA model in (4.4)–(4.6).